OpenStudy (anonymous):

IF YOU CAN DO THIS !!!! UR A GENIOUS!!! If KSD ~ WAT, what is the measure of w? Let k=16, d=34 and t=51

OpenStudy (marihelenh):

\[\frac{ x }{ 16 }=\frac{ 51 }{ 34 }\]

OpenStudy (marihelenh):

x=w

OpenStudy (anonymous):

25.4?

OpenStudy (anonymous):

Thank you guys!!!! Ur super helpful

OpenStudy (marihelenh):

No, it would just equal 24. \[(34)(w)=(16)(51)\]

OpenStudy (anonymous):

Yes I got 24 thank you

OpenStudy (anonymous):

Ill give u a medal too I just can't u both so damn helpful

OpenStudy (marihelenh):

?